Four forces, 60N at 60o and 50N at 30o both clockwise, 40N due north – west and 30N south – west, act at the origin. Calculate the magnitude and direction of the resultant force.


1
Expert's answer
2021-03-08T08:32:29-0500

"F_x=-40\cdot\cos45°-30\cdot\cos45°+50\cdot\sin30°+60\cdot\sin60°="


"=52\ (N)"


"F_y=40\cdot\sin45°-30\cdot\sin45°+50\cdot\cos30°+60\cdot\cos60°="


"=80.4\ (N)"


"F=\sqrt{52^2+80.4^2}=96\ (N)" . Answer


"\alpha=\tan^{-1}\frac{80.4}{57.1}=57.1°" North of East . Answer